40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 17136     Entry time: Thu Sep 8 12:01:02 2022
Author: JC 
Type: Configuration 
Category: Lab Organization 
Subject: Lab Organization 

The floor cable cover has been changed out for a new one. This is in Section X11.

Attachment 1: F41AD1DA-29E9-4449-99CB-5F43AE527CA6_1_105_c.jpeg  231 kB  Uploaded Thu Sep 8 13:04:14 2022  | Hide | Hide all
F41AD1DA-29E9-4449-99CB-5F43AE527CA6_1_105_c.jpeg
Attachment 2: FF5F2CE8-85E8-4B6F-8F8A-9045D978F670.jpeg  3.149 MB  Uploaded Thu Sep 8 14:18:44 2022  | Hide | Hide all
FF5F2CE8-85E8-4B6F-8F8A-9045D978F670.jpeg
ELOG V3.1.3-